aubreymathis:

A total of 17,100 seats are still available for the next hockey game. If 62% of the tickets are sold, how many seats are in the arena?Select the correct answer. A. 27,900 seats B. 45,000 seats C. 7,900 seats D. 7,100 seats

generation2006:

I think it might be B but i'm not sure.

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @generation2006 I think it might be B but i'm not sure. \(\color{#0cbb34}{\text{End of Quote}}\) how is it correct, u need to explain ur answer

aubreymathis:

idk but i have c for now

MAMASBREIA:

let me check

aubreymathis:

ok

aubreymathis:

I also need help on this question Roberta is thinking about joining the food co-op store. If she pays $10.00 to join, she will get 10% off on all the groceries she buys there. How much must she buy before her savings equal her membership fee? Select the correct answer. A. $10.00 B. $50.00 C. $100.00 D. $1,000.00

carmelle:

ok so,62% of the tickets were sold so that means 38% seats are still available. 38/100x = 17100 Next, we multiply both sides by 100 What do we get when we multiply both sides by 100?

carmelle:

\[\frac{ 38 }{ 100 }x \times 100 = ?\]

aubreymathis:

38

aubreymathis:

right?

carmelle:

correct

carmelle:

17,100 x 100 = ?

aubreymathis:

1710000?

carmelle:

yes, so now, \[38x =17,10000\]

carmelle:

Next we divide both sides by 38

carmelle:

what do we get when we do that?

aubreymathis:

x=0.45???

aubreymathis:

sorry LOL x=45,000

carmelle:

Yes!! \[x = 45,000\]

aubreymathis:

Thank you so much could you help me with anoter quetion

carmelle:

sure

carmelle:

post it and I'll help

aubreymathis:

on here?

carmelle:

no, just make another post

aubreymathis:

Roberta is thinking about joining the food co-op store. If she pays $10.00 to join, she will get 10% off on all the groceries she buys there. How much must she buy before her savings equal her membership fee? Select the correct answer. A. $10.00 B. $50.00 C. $100.00 D. $1,000.00

aubreymathis:

oop